For the balanced chemical reaction below:

P4O10 (g) + 6 H2O (l) → 4 H3PO4 (aq)

(a) Calculate the moles of water if there are 9.19 grams of tetraphosphorus decoxide.

(b) Calculate the grams of water if there are 0.730 moles of tetraphosphorus decoxide.

(c) Calculate the grams of tetraphosphorus decoxide if there are 5.41 grams of H3PO4.

(d) Calculate the moles of H3PO4 if there are 0.604 moles of tetraphosphorus decoxide.

I am really just unsure how to set up the formula.
Answer #1

a) molar mass of P4O10 = 283.889 gm/mol
    moles of P4O10 = 9.19 gms / 283.889 gm/mol = 0.032371 mol of P4O10
   moles of water = 6*0.032371 = 0.19423 mol of H2O
b) mol of P4O10 = 0.730 mol
   moles of water = 6 * 0.730 = 4.38 moles of H2O
   grams of water = 4.38 mol of H2O * 18 gm/mol = 78.84 gm of H2O
c) molar mass of H3PO4 = 98 gm/mol
   mol of H3PO4 = 5.41 gm / 98 gm/mol = 0.055204 mol
  moles of P4O10 required = (0.055204 mol / 4) *1 = 0.013801mol
  grams of P4O10 required = 0.013801mol * 283.889 gm/mol = 3.91795 gms of P4O10
d) moles of H3PO4 = 4*0.604 mol of P4O10 = 2.416 mol of H3PO4